OpenStudy (anonymous):

Find the greatest common factor: 24y8 + 6y6 (Points : 1) 3y2 6y6 6y8 24y8

OpenStudy (mathstudent55):

First look at the coefficients, 24 and 6. What is the GCF of 24 and 6?

OpenStudy (jdoe0001):

\(\bf 24y^8 + 6y^6\implies 4\cdot 6\cdot y^6\cdot y^2+6\cdot y^6\) any ideas?

OpenStudy (mathstudent55):